# Learning interactions between Rydberg atoms

### v1

@@ II.2 Bijection between interactions and spin-spin correlators
In this section, we prove a general result that provides a theoretical foundation for the Hamiltonian reconstruction approach implemented in this work. In part, we show that it suffices to measure NN spin correlation functions to determine uniquely the NN relative distances between atoms and, therefore, the full set of interactions. We show that there is, in general, a bijective correspondence relating the matrix of spatially dependent correlation functions $C_{i,j}$ to the matrix of interactions $J_{i,j}$ between Ising spins in the TFIM. This result can be viewed as a generalization of the Hohenberg-Kohn (HK) theorem of density functional theory [67, 68], which establishes a bijection between the local one-body potential and the local density of a system of interacting quantum particles. We note that previous work has demonstrated that a neural network can indeed be trained to learn the HK correspondence in interacting fermion models [69].
@@
First, we need two properties of the ground states of the Hamiltonian when $\Omega\neq 0$. We use |$\ket{\sigma}\equiv\ket{\sigma_{1},\cdots,\sigma_{N}}$ to denote the eigenstates of all the $\hat{\mathbf{S}}^{z}_{i}$’s.
@@
First, it is enough to prove the result for $\Omega<0$, since we can find a unitary transformation $U:=\prod_{j}(i\hat{\sigma}_{j}^{y})$ that changes the sign of $\Omega$ and $\delta$ but leaves $\{J_{ij}\}$ intact. Since each $i\hat{\sigma}_{j}^{y}$ locally maps $\ket{\uparrow}$ to $-\ket{\downarrow}$ and $\ket{\downarrow}$ to $\ket{\uparrow}$ at the $j$-th site, the |$\ket{\sigma}$’s are permuted among each other modulo an extra sign. Hence, the statement proved for the transformed Hamiltonian is true for the original sign.
Second, it is enough to prove that $V=e^{-\beta\hat{H}},\beta>0,$ is a positive matrix, namely it has all positive entries. Then the Perron-Frobenius theorem [70, 71, 72] tells us that the dominant (Perron) eigenvalue is simple (non-degenerate) and the corresponding (Perron) eigenvector has all positive components. Note that $V$ is Hermitian with positive eigenvalues with the Perron eigenvalue being $\exp(-\beta E)$ when $E$ is the ground-state energy of $\hat{H}$ and the Perron eigenvector is the ground state |$\ket{\Psi}$. Now we break up $\hat{H}=\hat{H}_{0}+\Omega\sum_{i}\hat{\sigma}_{i}^{x}$ where
@@
where $\hat{\sigma}_{i}^{x}(\tau):=e^{\tau\hat{H}_{0}}\hat{\sigma}_{i}^{x}e^{-\tau\hat{H}_{0}}$. This operator causes a spin flip at site $i$. This convergent series has every term that is nonnegative. Moreover between any |$\ket{\sigma}$ and |$\ket{\sigma^{\prime}}$, there is at least one term with at most $N$ $\hat{\sigma}_{i_{k}}^{x}$’s that is nonzero. Hence $V$ has all positive entries. ∎
Now we come to the main theorem.
Theorem (Generalization of the Hohenberg-Kohn theorem)
For the TFIM as defined in Eq. (2), for fixed non-zero $\Omega$ and fixed $\delta$, there exists a bijection between the correlation functions as $c_{i,j}=\langle\hat{\mathbf{S}}_{i}^{z}\hat{\mathbf{S}}_{j}^{z}\rangle$ and the set of interactions $J_{i,j}$, $J_{i,j}\Leftrightarrow c_{i,j}$.
@@
The direction $J_{i,j}\implies c_{i,j}$ is obvious. We now prove that $c_{i,j}\implies J_{i,j}$ by reductio ad absurdum, following the same logic for the classic HK theorem [67]. By the aforementioned lemma, the ground state of $\hat{H}$ is non-degenerate.
@@
To arrive at a contradiction, we will show that $J_{i,j}^{(1)}=J_{i,j}^{(2)}$ for every $i,j$ pair with $i<j$. After picking one such specific ${i,j}$ pair, choose all configurations where $\sigma_{i}=1,\sigma_{j}=1$. Call this set $s_{ij}$. When $i^{\prime}<j^{\prime}$, $\sum_{\sigma\in s_{ij}}\sigma_{i^{\prime}}\sigma_{j^{\prime}}=2^{N-2}\delta_{ii^{\prime}}\delta_{jj^{\prime}}$, $\delta_{ij}$ being the Kroenecker delta. Summing Eq. (8) over $\sigma\in s_{ij}$, we get $02^{N-2}(J_{i,j}^{(1)}-J_{i,j}^{(2)})=$ 0. This is in contradiction to the couplings being distinct. Hence, $\Psi_{1}\neq\Psi_{2}$.
@@
II.3 Machine learning techniques
As previously mentioned, the physical setup of interest consists of Rydberg atoms arranged on a 2D surface and maintained in place with optical tweezers. The measurement outcomes of this system, in terms of magnetization and spin-spin correlators, can be natively represented as a graph, see Fig. 2(a). For such a data representation, a well-suited ML architecture comes from the wide class of graph neural networks (GNN). The advantage that is especially important in the context of this work is that they preserve the graph structure of the processed data and can be made size invariant. Specifically, we use the Principal Neighbor Aggregation (PNA) variant with multiple aggregators [74]. We present some elements of the GNN in Fig. 2 and explain how it processes the physical data in Sec. II.3.2. To train the GNN, the $L^{2}$ norm between targets $\mathbf{y}$ and predictions $\hat{\mathbf{y}}$ is evaluated, where the targets and predictions correspond to relative nearest-neighbor (NN) distances ($R_{ij}\equiv R_{\langle ij\rangle}$) and relative next-nearest-neighbor (NNN) distances ($R_{ij}\equiv R_{\langle\langle ij\rangle\rangle}$) between atoms
[…]
II.3.1 Physical inputs
All the physical observables considered, namely the magnetization and spin-spin correlation functions, are computed using DMRG [77, 78]. There are two distinct ways to evaluate the observables. The first one, employed by default if not explicitly stated otherwise, is to measure directly the converged DMRG wave function. The second one is to sample $N_{\text{smpl}}$ times ($N_{\text{smpl}}$ snapshots) the wave function to compute the observables; it is efficient to ‘perfectly sample’ from Matrix Product States (MPS) [79, 80], so, assuming one has the converged ground state, one can draw perfect samples from it using the Born rule. For the two scenarios, we compute the observables in the Z-basis by default.
When computing the observables using the Born rule, one can sample the MPS expressed in the Z- or X-basis and then measure in the Z-basis to compute $\chi^{z}$ or $\chi^{x}$, respectively. Indeed, from Eq. (14),
$\chi^{x}_{i,j;\vec{\Omega}}=\sum_{\sigma}\left|\bra{\sigma}\hat{H}_{\text{Hd}}\ket{\Psi}_{\vec{\Omega}}\right|^{2}\bra{\sigma}\hat{\sigma}^{z}_{i}\hat{\sigma}^{z}_{j}\ket{\sigma},$ (15)
Again, in Eq. (15), the set {|$\{\ket{\sigma}\}$ denotes the eigenstates of the $\hat{\sigma}_{k}^{z}$’s.
In the scenario where we sample the MPS to compute observables, we resort to Eq. (15) to compute $\chi^{x}$. We combine the Z-observable ($\chi^{z}$) and X-observable ($\chi^{x}$) in the GNN training at some explicit occasions (case #$\#6$ in Table 1). Moreover, by default, if not explicitly stated otherwise – like in case #$\#6$ in Table 1 –, the inputted observables are Z-observables. The technical details related to the training and testing of the GNN are discussed further in Appendix A. As for the technical details of the DMRG calculations, these are summarized in Appendix B.
II.3.2 Graph neural network
The GNN layer’s input is a graph, and its output is a processed graph of the same structure but modified node (and possibly edge) values. In Fig. 2(a), we show an exemplary input graph that is relevant to this work and describes the outcomes of measurements on a three-spin chain as discussed in Sec. II.3.1. The GNN layer processes the input graph by sending learnable messages $m_{i\rightarrow j}$ between all pairs of directly connected nodes $i$ and $j$, as shown in Fig. 2(b). All messages within one GNN layer are outputs of the same neural network whose parameters $\mathbf{\theta}$ are optimized to send messages that are relevant to the task at hand. The input of this message-passing neural network is the values of the node pair between which the message is sent (here, $M_{i}$ and $M_{j}$) and the value of the edge connecting the node pair (here, $\chi_{ij}$). The same neural network with the same $\mathbf{\theta}$ can send different messages between various pairs of nodes because its input ($M_{i}$, $M_{j}$, and $\chi_{ij}$) varies between them. In our case, the message-passing network is a fully-connected neural network. The architectural details are in Appendix A.
We design the full neural network architecture to be independent of the size of the physical system. While the GNN layer natively has this property, ultimately, we also need a task network that makes predictions based on the hidden representation of the graph learned by the GNN layer(s). Therefore, we combine the GNN layers with a neural network parametrized by $\mathbf{\theta^{\prime\prime}}$ that takes as an input hidden values of only two NN nodes $h_{i}$ and $h_{j}$ along with their spin-spin correlation $\chi_{ij}$ and predicts the distance between NN spins $i$ and $j$, $R_{\langle ij\rangle}$, as presented in Fig. 2(c). The input size independence is achieved here in a similar way as in the GNN layer. One neural network learns a mapping between an NN pair and a distance between them and then makes predictions for all possible NN pairs. It can give different predictions for various NN pairs because the hidden values of participating nodes serving as the neural network input are different. In our case, the task network is a fully-connected neural network with two hidden layers. More generally, task networks can be adjusted to the form of Hamiltonian terms whose parameters they need to learn. For one-body Hamiltonian terms, the task network can take one $h_{i}$ as an input. For $n$-body terms, the input can include hidden values of $n$ nodes.
II.3.3 Training scenarios
We have devised several training scenarios in this work to compare different types of input data and architectures and select the best learning procedure. These training scenarios are listed in Table 1 and described in the following. To start with, i) the local magnetization $\langle\hat{\sigma}_{i}^{z}\rangle_{0;\vec{\Omega}}$, piled up across $\vec{\Omega}$, is inputted to the GNN without providing any physical information to the edges of the graph that stores the information about the cold atom array. Only the disorderless edge distances are provided as edge features. Then, ii) in addition to the local magnetization, the NN spin correlation functions (Eq. (14)) stacked up along $\vec{\Omega}$ are inputted to the GNN input layer to dress up the NN graph edges. This scenario should substantially improve the performance of the neural network according to the bijective relation between NN spin correlators $\chi^{\text{NN}}$ and spin couplings $J_{i,j}$ (see Sec. II.2 for the proof). Next, iii) we add the NNN spin correlation functions $\chi^{\text{NNN}}$ to the input data, thereby supplementing the NNN graph edges with physical information. In order to evaluate and discriminate the physical benefits from the architectural ones, we also consider scenarios where the features of NNN graph edges are set to unity to help the flow of information during optimization. This way, the NNN edges do not carry any information about $\chi^{\text{NNN}}$. Hence, iv) the features of the NNN graph edges are set to unity across $\vec{\Omega}$ to check out the effect of including NNN edges in the GNN architecture without assigning them any physical information (no values of $\chi^{\text{NNN}}$). The NNN edges per se could act as ‘skip connections’ in the back-propagation of the gradients. To continue the comparisons, in light of the bijection relation covered in Sec. II.2, v) we remove the physical content from the NNN graph edges (no values of $\chi^{\text{NNN}}$) and graph nodes (no values of local magnetization $M$), and set the features to unity across $\vec{\Omega}$. To finish the comparisons, vi) we combine the Z- and X-observable measurements of the NN and NNN correlators (Eq. (14)) in the training dataset. When computing X-observables by sampling the MPS, we resort to Eq. (15). The magnetization along $z$ is also considered in the training, and a $\Omega$-history is used. All the training scenarios are summarized in Table 1, and they all use the same $\Omega$-history (see further down below).
Case #$\#1$ $M$ and $\Omega$-history
Case #$\#2$ $M$, $\chi^{\text{NN}}$ and $\Omega$-history
Case #$\#3$ $M$, $\chi^{\text{NN}}$, $\chi^{\text{NNN}}$ and $\Omega$-history
Case #$\#4$ $M$, $\chi^{\text{NN}}$, NNN edges and $\Omega$-history
Case #$\#5$ $\chi^{\text{NN}}$, NNN edges and $\Omega$-history
Case #$\#6$ $M$, $\chi^{\text{NN}}$, $\chi^{\text{NNN}}$ and $\Omega$-history (Z+X)
To generate the graph dataset fed into the GNN, we choose a nominal NN distance between atomic sites $\bar{R}_{\langle i,j\rangle}$ of 10 $\mu$m, corresponding to the array spacing $a$ [5] – the notations $a$ and $\bar{R}_{\langle i,j\rangle}$ will be used interchangeably to denote the nominal NN distance. The positions of the Rydberg atoms are slightly shifted about their nominal values, with perturbations comprised within the bracket $\Delta S\in\left[-0.1,0.1\right]\ \mu$m, corresponding to the width of the white box shown in the phase diagram in Fig. 3. $\Delta S$ is drawn from a uniform distribution. Hence, the relative atomic distance $R_{i,j}=\bar{R}_{i,j}+\Delta R_{i,j}$, where the perturbation in the relative distance between atoms lies within $\Delta R_{i,j}\in\left[-0.2,0.2\right]\ \mu$m. The white box in Fig. 3 spans the allowed values of the transverse field $\Omega$ selected during the training phase.
Component-wise, the GNN edge-feature space is made of copies of the all-to-all exchange interactions $J_{i,j}$ at different equally-spaced values of $\Omega$. The explicit array of ten $\Omega$’s, used by all the scenarios laid out in Table 1, expands as such (the prefactor distributes across all array entries)
$\vec{\Omega}=100\cdot\left[-\frac{9}{9},-\frac{7}{9},\cdots,\frac{7}{9},\frac{9}{9}\right]\ \text{rad}\cdot\mu s^{-1},$
and is enclosed in the white box of Fig. 3.
II.3.4 Metrics
While investigating the scenarios listed above in Sec. II.3, we use three main metrics to evaluate performance. Throughout this work, the figure layout orders the metrics from top to bottom. Hence, we first plot in figures the coefficient of determination $R^{2}$, defined as
$R^{2}=1-\frac{\sum_{i}(y_{i}-\hat{y}_{i})^{2}}{\sum_{i}(y_{i}-\bar{y}_{i})^{2}},$ (16)
$\text{MAE}=\sum_{i=1}^{N}\frac{\left|\hat{y}_{i}-y_{i}\right|}{N},$ (17)
whose dimension is in nanometers. In Eq. (17), the sample size is denoted by $N$, and it grows larger as the cluster size increases since the number of graph edges and nodes increases as well. Note that the extrema of the possible variations (perturbations) of the NN relative distance is $\Delta R=\pm 0.2\ \mu$m. Then, at last, the median of the absolute error, denoted MEDAE for short, is plotted. The combined estimation of MEDAE and MAE allows us to evaluate the symmetry of the learned probability distribution: if it is skewed, the MEDAE and MAE differ significantly, whereas if it is symmetric, both the MAE and MEDAE look very similar.
III Results
As pointed out in Sec. II.3, there are two distinct ways to estimate the physical observables from the DMRG ground state. One way is to work out the expectation values having access to the full wave function content, while the other is to ‘sample’ the wave function thereby producing measurement snapshots. The former is useful to test thoroughly and faithfully the limitations and strengths of the GNN architecture employed, leaving out extra sources of statistical noises that would come from the projective measurements – only the numerical round-off error would be the source of ‘noise’. Therefore, the extent to which the bijection relation proven in Sec. II.2 is respected by the PNA GNN can be assessed. The performance of the GNN training using this ‘exact’ scheme is presented in Sec. III.2. The GNN training scheme utilizing snapshot-based observables, in closer relation to experiments, is presented in Sec. III.3. These two ways of measuring observables should coincide perfectly with each other in the limit where the number of snapshots tends to infinity. Before delving into GNN Hamiltonian learning results of Secs. III.2 and III.3, we ought to study more closely the phase diagram of the TFIM at the constant zero detuning field ($\delta=0$ in Eq. (1)). This is done in Sec. III.1.
III.1 Phase diagram
Figure 3: Phase diagram of the Rydberg TFIM for a $5\times 5$ square lattice. The white box encloses the parameter sets that can possibly be drawn during the training of the GNN. Along the $y$-axis, $a$ represents the nearest-neighbor distance between atoms. The inset plot zooms into the region of the phase diagram enclosed within the black box. Note that the phase diagram was calculated for an equally-spaced ordered lattice.
We then zoom in on the vicinity of the phase transition in Fig. 3 (inset plot) to see better the variation of the magnetization. The range of nominal NN distances $a$ corresponds to the height of the white box appearing in Fig. 3. Note that its height encompasses all the possible $a$’s given that $\Delta R_{i,j}\in\left[-0.2,0.2\right]\mu$m. In the region with a significant color gradient, the variance of the NN spin correlation functions is highest, as will be seen later in Fig. 6. $\Omega\gtrsim 10\,\text{rad}\cdot\mu s^{-1}$ corresponds to the point where the detuning and interaction terms are of the same order of magnitude, and thus, there is more variability in the ground state.
III.2 Predictions from the exact wave function
III.2.1 Discriminating between training scenarios
We illustrate in Fig. 4 the indicative metrics for all training scenarios covered in Sec. II.3 (see Table. 1), following through the figure layout described in Sec. II.3.4. The error bars show the standard error on the GNN metric results based on a set of five independently trained GNNs within each case scenario. From looking at Fig. 4, it seems like the physical information provided by the various correlations functions, namely the NNN spin correlation function and local magnetization, do not benefit as much from the training as the architectural add-ons like the NNN edges; it is even more obvious when the local magnetization is set to $1$ (case #$\#5$ in Table 1). Nonetheless, including the physical information coming from $\chi^{\text{NNN}}$ is not without benefits, as can be seen from comparing the case #$\#3$ (extra $\chi^{\text{NNN}}$) with #$\#2$ since the former case outperforms the latter as far as MAE (17) and MEDAE are concerned. By very far, the worst GNN training implementation is case #$\#1$, where absolutely no physical edge information – no NN nor NNN correlation function – is included in the graphs of the dataset. When adding the correlators measured in the X basis (Eq. (15)) like done in case #$\#6$, the prediction power of the model can be increased further, amassing more physical input through the training, at the cost of a higher standard error on the extrapolation of large cluster sizes.
Moreover, it looks like the additional physical information to $\chi^{\text{NN}}$ is detrimental to the predictions of the GNN for the cluster sizes part of the training dataset (before the grey vertical line); this can be seen in all three measures shown in Fig. 4. This extra information also seems to negatively impact the error for large cluster sizes: the smallest error is obtained using the case #$\#5$ while the largest is obtained using the case #$\#6$. The GNN is also more faithful to the bijection relation described in Sec. II.2 when no physical observables are supplemented to NNN graph edges or nodes. Indeed, compared to case #$\#5$, both case #$\#4$, corresponding to the addition of local magnetizations to nodes, and case #$\#3$, corresponding to the additions of local magnetizations to nodes and $\chi^{\text{NNN}}$ to NNN graph edges, diminish the performance.
Figure 5: $\Omega$-resolved performance of the training case #$\#3$. Top panel: $R^{2}$ as a function of Rydberg cluster sizes at various values of $\Omega$ (see legend). Middle panel: Mean absolute error of the predicted nearest-neighbor (NN) relative displacements. Bottom panel: Median of the mean absolute error of the relative NN displacements. $\Omega$ is shown in units of $\text{rad}\cdot\mu s^{-1}$.
III.2.2 Effect of $\vec{\Omega}$-histories
In the training scenarios covered in Fig. 4, we only considered cases where histories of $\Omega$ are used to enrich both the edge and node features in the graphs. Of course, one can wonder what happens if one selects solely one value of $\Omega$. For that matter, we show in Fig. 5 a collection of single $\Omega$ used to train the GNN following case #$\#3$, in that the physical observables like $M_{i}$, $\chi^{\text{NN}}$ and $\chi^{\text{NNN}}$ are used as node and edge attributes, but the dataset is limited to only one $\Omega$ instead of a $\Omega$-history. From looking at Fig. 5, it is striking to see how the training performance hinges on the selected value of $\Omega$. We nevertheless can confidently state that the values of $\Omega$ chosen above $\Omega\simeq C_{6}/\bar{R}^{6}_{\langle i,j\rangle}\sim 10\ \text{rad}\cdot\mu s^{-1}$, although not too far away, give significantly better results. Furthermore, above that threshold, in the inset plot of Fig. 3, for a fixed value of $\Omega$, the magnetization $M$ (18) varies more within the same set of $a$’s.
To better understand why some values of $\Omega$ perform better than others in Fig. 5, we plot in Fig. 6 the standard deviation (STD) of $\chi^{\text{NN}}$, denoted $\text{STD}[\chi^{\text{NN}}]$, as a function of the cluster sizes for all $\Omega$’s considered. In the range $\Omega\simeq\left[10,25\right]\ \text{rad}\cdot\mu s^{-1}$, fluctuations in the correlation function surpass that of all other $\Omega$’s by an order of magnitude at least, signaling the magnetic phase transition in the vicinity. Moreover, in that regime, the trend of $\text{STD}[\chi^{\text{NN}}]$ is far from being constant with the cluster size. Indeed, spin interactions in the critical regime become more long-range, which renders the training more size-dependent and greatly reduces extrapolation performance at larger cluster sizes – physics is different between cluster sizes. Even for $\Omega=25\ \text{rad}\cdot\mu s^{-1}$, the fluctuations increase for all values of cluster sizes, in line with the critical behavior of correlation functions when passing a quantum phase critical point. When the quantum critical point is passed, but the standard variation of $\chi^{\text{NN}}$ remains roughly within $1\times 10^{-2}\gtrsim\text{STD}[\chi^{\text{NN}}]\gtrsim 1\times 10^{-3}$, this is where we see an optimal gain in performance. Hence, when $\Omega>J_{\langle i,j\rangle}$, but not too much, one finds a sweet spot. However, one could generally take a $\Omega$ history for $\Omega>J_{\langle i,j\rangle}$. The values of $\Omega=\{0,5,300\}\ \text{rad}\cdot\mu s^{-1}$ all feature a $\text{STD}[\chi^{\text{NN}}]$ that falls below $1\times 10^{-3}$, too small for the GNN to pick up useful signals.
Figure 7: Training dataset cluster size effect. Incidence of the cluster sizes included in the training dataset on various metrics. In blue, clusters of sizes $4\times 4$ and $5\times 5$ were considered, whereas in orange, one extra cluster size $6\times 6$ was considered. The vertical lines refer to the point beyond which sizes lie outside of the training range, i.e. these sizes did not take part of the training datasets. The colors of the vertical lines match with that of the curves.

### v2

@@ Appendix A Proof of bijection between interactions and spin-spin correlators
In this section of the Appendix, we prove a general result that provides a theoretical foundation for the Hamiltonian reconstruction approach implemented in this work. In part, we show that it suffices to measure NN spin correlation functions to determine uniquely the NN relative distances between atoms and, therefore, the full set of interactions. We show that there is, in general, a bijective correspondence relating the matrix of spatially dependent correlation functions $c_{i,j}$ to the matrix of interactions $J_{i,j}$ between Ising spins in the TFIM. This result can be viewed as a generalization of the Hohenberg-Kohn (HK) theorem of density functional theory [72, 73], which establishes a bijection between the local one-body potential and the local density of a system of interacting quantum particles, and as a quantum many-body generalization of the Henderson theorem [74]. We note that previous work has demonstrated that a neural network can indeed be trained to learn the HK correspondence in interacting fermion models [97].
@@
where $i<j$ indicates that each link is counted only once. We define the correlation functions as: $c_{i,j}=\langle\hat{\mathbf{S}}_{i}^{z}\hat{\mathbf{S}}_{j}^{z}\rangle$, where $\langle\cdot\rangle$ indicates the expectation value and $\hat{\mathbf{S}}_{i}^{z}$ is the spin operator in the Z-basis acting on the $i$-th site, related to the Z Pauli matrix via $\hat{\mathbf{S}}_{i}^{z}=\frac{\hbar}{2}\mathbf{\hat{\sigma}}_{i}^{z}$. We prove that, for fixed $\Omega\neq 0$ and $\delta$, there is a unique set of interactions $J_{i,j}$ that yields a set $c_{i,j}$, and conversely: $J_{i,j}\Leftrightarrow c_{i,j}$. The condition $\Omega\neq 0$ is important since, for example, when $\Omega=\delta=0$, for all choices of $\{J_{i,j}\}$ with each individual $J_{i,j}<0$ the ground-state is the fully polarized ferromagnet and the $c_{i,j}$’s are obviously the same, irrespective of the specific values of the interactions.
@@
[…]
@@
$=\sum_{n=0}^{\infty}(-\Omega)^{n}$ $\sum_{i_{1},i_{2}\cdots,i_{n}}\int_{\tau_{n-1}}^{\beta}d\tau_{n}\cdots\int_{\tau_{1}}^{\beta}d\tau_{2}\int_{0}^{\beta}d\tau_{n}$
@@
Now we come to the main theorem, adapted to the TFIM, which assumes the $J$-representability of the connected correlation functions [70, 71].
Theorem (Hohenberg-Kohn-Henderson Theorem for the TFIM)
For the TFIM as defined in Eq. (8), for fixed nonzero $\Omega$ and fixed $\delta$, there exists a bijection between the $J$-representable correlation functions $c_{i,j}=\langle\hat{\mathbf{S}}_{i}^{z}\hat{\mathbf{S}}_{j}^{z}\rangle$ and the set of interactions $J_{i,j}$, $J_{i,j}\Leftrightarrow c_{i,j}$.
@@
The direction $J_{i,j}\implies c_{i,j}$ is obvious. We now prove that $c_{i,j}\implies J_{i,j}$ by reductio ad absurdum, following the same logic for the classic HK theorem [72]. By the aforementioned lemma, the ground state of $\hat{H}$ is nondegenerate.
@@
To arrive at a contradiction, we will show that $J_{i,j}^{(1)}=J_{i,j}^{(2)}$ for every $i,j$ pair with $i<j$. After picking one such specific i,j{i,j} pair, choose all configurations where $\sigma_{i}=1,\sigma_{j}=1$. Call this set $s_{ij}$. When $i^{\prime}<j^{\prime}$, $\sum_{\sigma\in s_{ij}}\sigma_{i^{\prime}}\sigma_{j^{\prime}}=2^{N-2}\delta_{ii^{\prime}}\delta_{jj^{\prime}}$, $\delta_{ij}$ being the Kroenecker delta. Summing Eq. (14) over $\sigma\in s_{ij}$, we get $2^{N-2}(J_{i,j}^{(1)}-J_{i,j}^{(2)})=0$. This is in contradiction to the couplings being distinct. Hence, $\Psi_{1}\neq\Psi_{2}$.
@@
Appendix B Number of physical system sizes in training datasets